OpenStudy (anonymous):

limit (sqrt{x^2 + 1} - ({x^2 + 2 x}/{x})) x-->2 find the limit

OpenStudy (zzr0ck3r):

\[\lim_{x\rightarrow 2}\sqrt{x^2+1}-\frac{x^2+2x}{x}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if the limit exists, lim x-> 2 of f(x) = f(2)

OpenStudy (zzr0ck3r):

just plug in 2, we are well behaved at x = 2

OpenStudy (anonymous):

 is this the answer